the length of a rectangle is 15m more than the width. the perimeter of the rectangle is 74m
Answer by waynest(281)

P = 2L + 2W
L = w + 15
74 = 2(w + 15) + 2w
74 = 2w + 30 + 2w
74 = 4w + 30
74 - 30 = 4w
44 = 4w
44 / 4 = w
11 = w